# Random Feature Representation Boosting

### v3

@@ 1.2 Related literature
Random feature models: Random feature models use fixed, randomly generated features to map data into a high dimensional space, enabling efficient linear learning. These methods encompass a broad class of models studied under various names, including random Fourier features (Rahimi & Recht, 2007; Sriperumbudur & Szabo, 2015; Li et al., 2019; Kammonen et al., 2022; Davis et al., 2024), extreme learning machines (Huang et al., 2004, 2012; Huang, 2014), random features or RFNNs (Huang et al., 2006; Rahimi & Recht, 2008a, b; Rudi & Rosasco, 2017; Carratino et al., 2018; Yehudai & Shamir, 2019; Mei & Montanari, 2022; Bolager et al., 2023; Lanthaler & Nelsen, 2023; Ayme et al., 2024), reservoir computing (Jaeger, 2001; Lukoševičius & Jaeger, 2009; Gallicchio et al., 2017; Grigoryeva & Ortega, 2018; Tanaka et al., 2019; Hart et al., 2020; Gonon et al., 2023, 2024), kernel methods (Kar & Karnick, 2012; Sinha & Duchi, 2016; Sun et al., 2018; Szabo & Sriperumbudur, 2019; Cheng et al., 2023; Wang et al., 2024; Wang & Feng, 2024), and scattering networks (Bruna & Mallat, 2013; Cotter & Kingsbury, 2017; Oyallon et al., 2019; Trockman et al., 2023). Recently, RFNNs and related approaches have demonstrated exceptional performance in both speed and accuracy across a wide variety of applications. These include solving partial differential equations (Nelsen & Stuart, 2021; Gonon, 2023; Gattiglio et al., 2024; Neufeld & Schmocker, 2024), online learning (Prabhu et al., 2024), time series classification (Dempster et al., 2020, 2023), and in mathematical finance (Jacquier & Zuric, 2023; Herrera et al., 2024). RFNNs have also been explored in the context of quantum computing (Innocenti et al., 2023; Gonon & Jacquier, 2023; Martínez-Peña & Ortega, 2023; Xiong et al., 2024), and in relation to random neural controlled differential equations and state-space models (Cuchiero et al., 2021; Cirone et al., 2023, 2024).
@@ 2.1 Traditional Gradient Boosting
To minimize the risk based on this approximation, a new weak learner $g_{t+1}$ is fit to the negative functional gradient, $-\nabla\mathcal{R}(F_{t})$, and the ensemble is then updated: $F_{t+1}=F_{t}+\eta\alpha_{t+1}g_{t+1}$, where $\eta,\alpha_{t+1}>0$ are the global and local learning rates. The functional gradient can be shown to be equal to
@@
and is easily computed for empirical measures $\widehat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ using the empirical risk ^$\widehat{\mathcal{R}}(F)$. The method for fitting $g_{t+1}$ varies between implementations and depends on the class of weak learners, as seen in modern approaches like XGBoost (Chen & Guestrin, 2016) which incorporates second-order information.
@@ 2.2 Gradient Representation Boosting
Unlike classical boosting, which boosts in label space, neural network gradient representation boosting aims to additively learn a feature representation by modelling $F$ via $F=F_{t}(X)=W_{t}^{\top}\Phi_{t}(X)$, where $\Phi_{t}=\sum_{s=0}^{t}\eta g_{s}$ is a gradient boosted feature representation and $W_{t}\in\mathbb{R}^{D\times d}$ is the top-level linear predictor. In other words, gradient representation boosting seeks to learn the feature representation $\Phi_{t}$ additively to build a single strong predictor, rather than constructing $F_{t}$ as an ensemble of weak predictors (Huang et al., 2018; Nitanda & Suzuki, 2018, 2020; Suggala et al., 2020). For a depth $t$ ResNet $\Phi_{t}$, defined recursively as
@@
Let ^$\widehat{\mathcal{R}}(W,\Phi)$ be the empirical risk, $\mathcal{W}$ the hypothesis set of top-level linear predictors, and $\mathcal{G}_{t}$ the set of weak feature transformations. As outlined by Suggala et al. (2020), there are two main strategies for training ResNets via gradient representation boosting, which we describe below.
@@
Potential Challenges: The gradient-greedy approach for constructing ResNets has remained relatively unexplored in the literature. Suggala et al. (2020), who introduced both strategies in the context of end-to-end trained networks, only mention in passing that the gradient-greedy strategy performed worse comparatively. We believe this might be attributed to the functional direction of $g$ failing to be properly preserved during training with SGD. Specifically, when $g$ is chosen from the family $g=h(\Phi_{t-1})$, where $h$ is a shallow neural network, minimizing $\langle g,\nabla_{2}\widehat{\mathcal{R}}(W_{t-1},\Phi_{t-1})\rangle_{L_{2}^{D}(\widehat{\mu})}$ via SGD might increase $\|g\|_{L_{2}^{D}(\widehat{\mu})}$ without ensuring that $g$ aligns with the functional gradient in $L_{2}^{D}(\widehat{\mu})$. Since the first-order approximation (4) only holds for $g$ with small functional norm, we argue that this objective should instead be minimized under the constraint $\|g\|_{L_{2}^{D}(\widehat{\mu})}=1$. In our random feature setting, we incorporate this constraint and show in Section 4 that it leads to the gradient-greedy approach outperforming the exact-greedy strategy.
@@ 3.2 Exact-Greedy Case for Mean Squared Error Loss
Recall that in the exact-greedy layer-wise paradigm, the objective at layer $t$ is to find a function $g_{t}\in\mathcal{G}_{t}$ that additively and greedily minimizes the empirical risk ^$\widehat{\mathcal{R}}(W_{t},\Phi_{t-1}+g_{t})$ for some linear map $W_{t}\in\mathcal{W}$, given a ResNet feature representation $\Phi_{t-1}$. We propose using functional gradient boosting to construct a residual block of the form $g_{t}=A_{t}f_{t}$, where $A_{t}\in\mathbb{R}^{D\times p}$ is a linear map, and $f_{t}$ is a random feature layer. We consider the cases where $A_{t}$ is a scalar multiple of the identity, a diagonal matrix, or a general dense matrix. The procedure is outlined below:
@@
where $r_{i}=y_{i}-W_{t-1}^{\top}\Phi_{t-1}(x_{i})$ are the residuals of the model at layer $t-1$. We term problems of the form (3.2) sandwiched least squares problems, for which Theorem 3.1 below provides the existence of closed-form analytical solutions which are fast to compute.
@@ 3.3 Gradient Boosting Random Features
which holds for functions $g$ with small $L_{2}^{D}(\mu)$-norm. As discussed in the context of general gradient representation boosting, a potential issue with directly minimizing $\langle g,\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)\rangle_{L^{D}_{2}(\widehat{\mu})}$ is that $g$ might learn to maximize its magnitude without following the direction of the functional gradient. To address this, we constrain the problem to ensure that $g$ maintains a unit norm in $L_{2}^{D}(\widehat{\mu})$. If we restrict $g$ to residual blocks of the form $g=Af$, where $A\in\mathbb{R}^{D\times p}$ is a linear map, and $f\in\mathbb{R}^{p}$ are random features, then solving the constrained $L_{2}^{D}(\widehat{\mu})$-inner product minimization problem becomes equivalent to solving a quadratically constrained least squares problem. See Appendix B for the proof.
@@
minimize $\textnormal{minimize }\quad\frac{1}{n}\sum_{i=1}^{n}\|\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)(x_{i})-Af(x_{i})\|^{2},$
subject to $\textnormal{subject to }\quad\frac{1}{n}\sum_{i=1}^{n}\|Af(x_{i})\|^{2}=1,$
@@
where $s$ is the softmax function, and $e_{y_{i}}$ is the one-hot vector for label $y_{i}$. For full derivation, see Appendix C. We then fit $A_{t}$ by fitting $f_{t+1}$ to the negative normalized functional gradient $-\frac{\sqrt{n}G}{\|G\|_{F}}$ via multiple least squares, according to Theorem B.1. In practice, we also include $\ell_{2}$ regularization.
@@ 3.4 Theoretical Guarantees
$+C\left(\frac{1}{T^{a}}+\frac{T^{2-s}\big{(}1+\sqrt{\log{\frac{T}{\delta}}}\big{)}}{\sqrt{\widetilde{n}}}\right),$
@@ 4.2 Point Cloud Separation
We evaluate RFRBoost on a challenging synthetic dataset originating from the neural ODE literature (Sander et al., 2021). The dataset consists of 10,000 points sampled from concentric circles, and the task is to linearly separate (i.e. classify) the concentric circles while restricting the ResNet hidden size to 2, see Figure 4. We compare RFRBoost to E2E MLP ResNets, while also presenting classification results for logistic regression and RFNNs in Table 3. All models were trained with cross entropy loss, using a 5-fold CV grid search for hyperparameter tuning. For E2E MLP ResNets, the learning rate was tuned, while for the other models, only the $\ell_{2}$ regularization of the classification head was tuned. The hidden size, residual block feature dimension, and activation function was fixed at $2$, $512$, and $\tanh$, respectively, for all models. See Section E.3 for more details.
@@ 5 Conclusion
This paper introduced RFRBoost, a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost addresses the limitations of single-layer RFNNs by using random features to learn a feature representation that approximates the functional gradient of the network, thereby enhancing performance while retaining the computational benefits of convex optimization for RFNNs. We derived closed-form solutions for greedy layer-wise boosting with MSE loss, and presented a general framework for arbitrary loss functions based on solving a quadratically constrained least square problem. Through extensive numerical experiments on tabular datasets for both regression and classification, we have demonstrated that RFRBoost significantly outperforms traditional RFNNs and end-to-end trained MLP ResNets in the small- to medium-scale regime where RFNNs are typically applied, while offering substantial computational advantages, and theoretical guarantees stemming from boosting theory. RFRBoost represents a significant step towards building powerful, stable, efficient, and theoretically sound deep networks using untrained random features. Future work will focus on extending RFRBoost to other domains such as time series or image data, exploring different types of random features and momentum strategies, implementing more efficient GPU acceleration, scaling to large datasets, and using RFRBoost as an initialization strategy for large-scale end-to-end training.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Analytic Solutions to Sandwiched Least Squares
@@
00 0 $=\frac{\partial J(a_{1},...,a_{D})}{\partial a_{k}}=\frac{1}{n}\sum_{i=1}^{n}-2\big{(}W_{k}X_{i,k}\big{)}^{\top}\big{(}R_{i}-W^{\top}[a_{1}X_{i,1},...,a_{D}X_{i,D}]^{\top}\big{)}+2\lambda a_{k}$
@@ Appendix B Functional Gradient Inner Product
where we see that the first term $\|H_{i}\|^{2}$ is constant w.r.t $A$, and the third term gives the constraint. Hence, minimizing the constrained least squares problem is equivalent to maximizing the inner product, and the solution to the original problem is obtained by multiplying the least squares solution by $-1$ since we are interested in the argmin rather than the argmax.
@@ D.1 Preliminaries
where $\rho_{i,j}$ are Rademacher random variables, that is, independent random variables taking values $1$ and $-1$ with equal probability.
@@
=$\mathcal{H}=\{x\mapsto A\tanh(Bx):\lambda_{\textnormal{max}}(A)\textnormal{ and }\lambda_{\textnormal{max}}(B)\textnormal{ bounded by }\lambda\}.$
@@
=$\mathcal{F}=\left\{x\mapsto\sum_{j=1}^{m}v_{j}\sigma(f_{j}(x)):f_{j}\in\mathcal{F}_{j},\|v\|_{1}\leq C,v\in\mathbb{R}^{m},\|v\|_{\infty}\leq A\right\}$
@@ D.2 RFRBoost Rademacher Computations
$\mathfrak{R}(\mathcal{G}_{t})=\mathcal{O}\left(\frac{pD^{\frac{3}{2}}\log(D)t^{1-s}}{\sqrt{\widetilde{n}}}\right).$
@@
$\leq 2\lambda_{1}D\sum_{j=1}^{p}\mathbb{E}_{\rho}\bigg{[}\sup_{\begin{subarray}{c}B\in\mathbb{R}^{p\times D}\\ \lambda_{\textnormal{max}}(B)\leq\lambda_{1}\end{subarray}}\frac{1}{\widetilde{n}}\sum_{i=1}^{\widetilde{n}}\rho_{i,1}\left\langle B_{j},\Phi_{t-1}(x_{t,i})\right\rangle\bigg{]}+D\mathcal{O}\left(\frac{\sqrt{p}\log p}{\sqrt{\widetilde{n}}}\right)$
$\leq 2\lambda_{1}^{2}pD\sqrt{D}\mathbb{E}_{\rho}\bigg{[}\frac{1}{\widetilde{n}}\left\|\sum_{i=1}^{\widetilde{n}}\rho_{i,1}\Phi_{t-1}(x_{t,i})\right\|_{\infty}\bigg{]}+D\mathcal{O}\left(\frac{\sqrt{p}\log p}{\sqrt{\widetilde{n}}}\right)$
$\leq\frac{4\lambda_{1}^{2}pD\sqrt{D}\log D}{\sqrt{\widetilde{n}}}\left\|\Phi_{t-1}(x_{t,i})\right\|_{\infty}+D\mathcal{O}\left(\frac{\sqrt{p}\log p}{\sqrt{\widetilde{n}}}\right)$
$\leq\frac{4\lambda_{1}^{3}pD\sqrt{D}\log(D)ct^{1-s}}{(1-s)\sqrt{\widetilde{n}}}+D\mathcal{O}\left(\frac{\sqrt{p}\log(p)}{\sqrt{\widetilde{n}}}\right)$
$=\mathcal{O}\left(\frac{pD^{\frac{3}{2}}\log(D)t^{1-s}}{\sqrt{\widetilde{n}}}\right).$
@@
$\mathfrak{R}(\mathcal{W},\mathcal{G}_{t})=\mathcal{O}\left(\frac{pdD^{\frac{3}{2}}\log(D)t^{1-s}}{\sqrt{\widetilde{n}}}\right).$
@@
$\leq^{(lemma)}2\lambda_{1}d\sum_{k=1}^{D}\mathbb{E}_{\rho}\bigg{[}\sup_{\begin{subarray}{c}g\in\mathcal{G}_{t}\end{subarray}}\sum_{i=1}^{\widetilde{n}}\rho_{i,1}g_{k}(x_{t,i})\bigg{]}+d\mathcal{O}\left(\frac{\lambda_{1}^{2}\sqrt{pD}\log D}{\sqrt{\widetilde{n}}}\right)$
$\leq^{(prev\>result)}2\lambda_{1}d\mathcal{O}\left(\frac{pD^{\frac{3}{2}}\log(D)t^{1-s}}{\sqrt{\widetilde{n}}}\right)+d\mathcal{O}\left(\frac{\lambda_{1}^{2}\sqrt{pD}\log D}{\sqrt{\widetilde{n}}}\right)$
$=\mathcal{O}\left(\frac{pdD^{\frac{3}{2}}\log(D)t^{1-s}}{\sqrt{\widetilde{n}}}\right).$

### v4

@@ 1.2 Related literature
Random feature models: Random feature models use fixed, randomly generated features to map data into a high dimensional space, enabling efficient linear learning. These methods encompass a broad class of models studied under various names, including random Fourier features (Rahimi & Recht, 2007; Sriperumbudur & Szabo, 2015; Li et al., 2019; Kammonen et al., 2022; Davis et al., 2024), extreme learning machines (Huang et al., 2004, 2012; Huang, 2014), random features or RFNNs (Huang et al., 2006; Rahimi & Recht, 2008a, b; Rudi & Rosasco, 2017; Carratino et al., 2018; Yehudai & Shamir, 2019; Mei & Montanari, 2022; Bolager et al., 2023; Lanthaler & Nelsen, 2023; Ayme et al., 2024), reservoir computing (Jaeger, 2001; Lukoševičius & Jaeger, 2009; Gallicchio et al., 2017; Grigoryeva & Ortega, 2018; Tanaka et al., 2019; Hart et al., 2020; Gonon et al., 2023, 2024), kernel methods (Kar & Karnick, 2012; Sinha & Duchi, 2016; Sun et al., 2018; Szabo & Sriperumbudur, 2019; Cheng et al., 2023; Wang et al., 2024; Wang & Feng, 2024), and scattering networks (Bruna & Mallat, 2013; Cotter & Kingsbury, 2017; Oyallon et al., 2019; Trockman et al., 2023). Recently, RFNNs and related approaches have demonstrated exceptional performance in both speed and accuracy across a wide variety of applications. These include solving partial differential equations (Nelsen & Stuart, 2021; Gonon, 2023; Gattiglio et al., 2024; Neufeld & Schmocker, 2024), online learning (Prabhu et al., 2024), time series classification (Dempster et al., 2020, 2023), and in mathematical finance (Jacquier & Zuric, 2023; Herrera et al., 2024). RFNNs have also been explored in the context of quantum computing (Innocenti et al., 2023; Gonon & Jacquier, 2023; Martínez-Peña & Ortega, 2023; Xiong et al., 2024), and in relation to random neural controlled differential equations and state-space models (Cuchiero et al., 2021; Cirone et al., 2023, 2024; Biagini et al., 2024).
@@ 2.1 Traditional Gradient Boosting
To minimize the risk based on this approximation, a new weak learner $g_{t+1}$ is fit to the negative functional gradient, $\nabla\mathcal{R}(F_{t})$, and the ensemble is then updated: $F_{t+1}=F_{t}+\eta\alpha_{t+1}g_{t+1}$, where $\eta,\alpha_{t+1}>0$ are the global and local learning rates. The functional gradient can be shown to be equal to
@@
and is easily computed for empirical measures $\widehat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ using the empirical risk $\widehat{\mathcal{R}}(F)$. The method for fitting $g_{t+1}$ varies between implementations and depends on the class of weak learners, as seen in modern approaches like XGBoost (Chen & Guestrin, 2016) which incorporates second-order information.
@@ 2.2 Gradient Representation Boosting
Unlike classical boosting, which boosts in target space, neural network gradient representation boosting aims to additively learn a feature representation by modelling $F$ via $F=F_{t}(X)=W_{t}^{\top}\Phi_{t}(X)$, where $\Phi_{t}=\sum_{s=0}^{t}\eta g_{s}$ is a gradient boosted feature representation and $W_{t}\in\mathbb{R}^{D\times d}$ is the top-level linear predictor. In other words, gradient representation boosting seeks to learn the feature representation $\Phi_{t}$ additively to build a single strong predictor, rather than constructing $F_{t}$ as an ensemble of weak predictors (Huang et al., 2018; Nitanda & Suzuki, 2018, 2020; Suggala et al., 2020). For a depth $t$ ResNet $\Phi_{t}$, defined recursively as
@@
Let $\widehat{\mathcal{R}}(W,\Phi)$ be the empirical risk, $\mathcal{W}$ the hypothesis set of top-level linear predictors, and $\mathcal{G}_{t}$ the set of weak feature transformations. As outlined by Suggala et al. (2020), there are two main strategies for training ResNets via gradient representation boosting, which we describe below.
@@
[…]
@@ 3.2 Exact-Greedy Case for Mean Squared Error Loss
Recall that in the exact-greedy layer-wise paradigm, the objective at layer $t$ is to find a function $g_{t}\in\mathcal{G}_{t}$ that additively and greedily minimizes the empirical risk $\widehat{\mathcal{R}}(W_{t},\Phi_{t-1}+g_{t})$ for some linear map $W_{t}\in\mathcal{W}$, given a ResNet feature representation $\Phi_{t-1}$. We propose using functional gradient boosting to construct a residual block of the form $g_{t}=A_{t}f_{t}$, where $A_{t}\in\mathbb{R}^{D\times p}$ is a linear map, and $f_{t}$ is a random feature layer. We consider the cases where $A_{t}$ is a scalar multiple of the identity, a diagonal matrix, or a general dense matrix. The procedure is outlined below:
@@
where $r_{i}=y_{i}-W_{t-1}^{\top}\Phi_{t-1}(x_{i})$ are the residuals of the model at layer t1t-1. We term problems of the form (3.2) sandwiched least squares problems, for which Theorem 3.1 below provides the existence of closed-form analytical solutions which are fast to compute.
@@ 3.3 Gradient Boosting Random Features
[…]
@@
$\textnormal{minimize }\quad\frac{1}{n}\sum_{i=1}^{n}\|\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)(x_{i})-Af(x_{i})\|^{2},$
$\textnormal{subject to }\quad\frac{1}{n}\sum_{i=1}^{n}\|Af(x_{i})\|^{2}=1,$
@@
where $s$ is the softmax function, and $e_{y_{i}}$ is the one-hot vector for label $y_{i}$. For full derivation, see Appendix C. We then fit $A_{t}$ by fitting $f_{t+1}$ to the negative normalized functional gradient $\frac{\sqrt{n}G}{\|G\|_{F}}$ via multiple least squares, according to Theorem B.1. In practice, we also include $\ell_{2}$ regularization.
@@ 3.4 Theoretical Guarantees
[…]
@@ 4.2 Point Cloud Separation
We evaluate RFRBoost on a challenging synthetic dataset originating from the neural ODE literature (Sander et al., 2021). The dataset consists of 10,000 points sampled from concentric circles, and the task is to linearly separate (i.e. classify) the concentric circles while restricting the ResNet hidden size to 2, see Figure 4. We compare RFRBoost to E2E MLP ResNets, while also presenting classification results for logistic regression and RFNNs in Table 3. All models were trained with cross entropy loss, using a 5-fold CV grid search for hyperparameter tuning. For E2E MLP ResNets, the learning rate was tuned, while for the other models, only the $\ell_{2}$ regularization of the classification head was tuned. The hidden size, residual block feature dimension, and activation function was fixed at $2$, 512512, and $\tanh$, respectively, for all models. See Section E.3 for more details.
@@ 5 Conclusion
This paper introduced RFRBoost, a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost addresses the limitations of single-layer RFNNs by using random features to learn optimal ResNet-like residual blocks that approximate the negative functional gradient at each layer of the network, thereby enhancing performance while retaining the computational benefits of convex optimization for RFNNs. This procedure can be viewed as performing functional gradient descent on the network neurons, which has connections to gradient boosting theory, as opposed to classical SGD which performs gradient descent on the network weights and biases. In our framework, we derived closed-form solutions for greedy layer-wise boosting with MSE loss, and presented a general fitting algorithm for arbitrary loss functions based on solving a quadratically constrained least squares problem. Through extensive numerical experiments on tabular datasets for both regression and classification, we demonstrated that RFRBoost significantly outperforms traditional RFNNs and end-to-end trained MLP ResNets in the small- to medium-scale regime where RFNNs are typically applied, while offering substantial computational advantages, and theoretical guarantees stemming from boosting theory. RFRBoost represents a significant step towards building powerful, stable, efficient, and theoretically sound deep networks using untrained random features. Future work will focus on extending RFRBoost to other domains such as time series or image data, exploring different types of random features and momentum strategies, implementing more efficient GPU acceleration, scaling to large datasets, and using RFRBoost as an initialization strategy for large-scale end-to-end training.
@@ Acknowledgements
Appendix A Analytic Solutions to Sandwiched Least Squares Problems
@@
$0$ $=\frac{\partial J(a_{1},...,a_{D})}{\partial a_{k}}=\frac{1}{n}\sum_{i=1}^{n}-2\big{(}W_{k}X_{i,k}\big{)}^{\top}\big{(}R_{i}-W^{\top}[a_{1}X_{i,1},...,a_{D}X_{i,D}]^{\top}\big{)}+2\lambda a_{k}$
@@ Appendix B Functional Gradient Inner Product
where we see that the first term $\|H_{i}\|^{2}$ is constant w.r.t $A$, and the third term gives the constraint. Hence, minimizing the constrained least squares problem is equivalent to maximizing the inner product, and the solution to the original problem is obtained by multiplying the least squares solution by 1-1 since we are interested in the argmin rather than the argmax.
@@ D.1 Preliminaries
where $\rho_{i,j}$ are Rademacher random variables, that is, independent random variables taking values $1$ and 1-1 with equal probability.
@@
$\mathcal{H}=\{x\mapsto A\tanh(Bx):\lambda_{\textnormal{max}}(A)\textnormal{ and }\lambda_{\textnormal{max}}(B)\textnormal{ bounded by }\lambda\}.$
@@
$\mathcal{F}=\left\{x\mapsto\sum_{j=1}^{m}v_{j}\sigma(f_{j}(x)):f_{j}\in\mathcal{F}_{j},\|v\|_{1}\leq C,v\in\mathbb{R}^{m},\|v\|_{\infty}\leq A\right\}$
@@ D.2 RFRBoost Rademacher Computations
[…]
@@
[…]
@@
[…]
@@
[…]